OpenStudy (anonymous):

let a>0. Show that the maximum value of h(x) is (2+a)/(1+a) h(x) = 1/(1+abs(x)) + 1/(1+abs(x-a))

OpenStudy (anonymous):

x>a\[h(x)=\frac{1}{1+x}+\frac{1}{1+x-a}\]0<x<a\[h(x)=\frac{1}{1+x}+\frac{1}{1+a-x}\]x<0\[h(x)=\frac{1}{1-x}+\frac{1}{1-x-a}\]

OpenStudy (anonymous):

and i have a question @Stamished are u allowed to use calculus?

OpenStudy (anonymous):

yes, definitely. I'd actually prefer if I could get an answer with Calculus (:

OpenStudy (anonymous):

thanks so much for helping

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

and dont forget to check bounds... x=0 and x=a

OpenStudy (anonymous):

hmmm..., I'm still confused on what I should do next...

OpenStudy (anonymous):

differentiate on each interval and find critical points...

OpenStudy (anonymous):

someone feel free to offer a better approach.

OpenStudy (anonymous):

kay, I'm going to try it now

OpenStudy (anonymous):

for example \(x\ge a\) \[h'(x)=-\frac{1}{(1+x)^2}-\frac{1}{(1+x-a)^2}<0\]decreasing for all x>a so the maximum of this interval take places at x=a u do the rest :)

OpenStudy (anonymous):

okay cool, so I'm doing it right so far. Thanks!! :D